Question 10 Tries remaining: 2 quantity without a tax is 330 units. The government levies a $85 per unit tax on the suppliers of sugar. Points out of 7.70 Calculate deadweight loss from this tax. The demand for sugar is given by: aD 420-0.25P. The supply of sugar is given by: Qs- 4P-1110. The equilibrium Flag question(Do not include a S sign in your response. Round to the nearest two decimal places if necessary.) Answer: Check
Answer #1

The quantity demanded (QD ) = 420-0.25P and quantity supplied (QS ) = 4P-1110
Under market equilibrium, QD = QS

i.e., 420-0.25P =4P-1110

or, 4.25P = 1530, or P = 360 and equilibrium quantity = 420-0.25*360= 330

Again if the government levies a tax of $85 per unit on suppliers of sugar, then new supplier function is

QS = 4P-85-1110

Under equilibrium market condition,

    QD = QS

i.e., 420-0.25P = 4P-85-1110

   4.25P = 1615 or, P = 380 and equilibrium quantity with tax = 4*360-85-1110= 245

Therefore, Dead-weight loss = 1/2 * difference in equilibrium prices * differences in equilibrium quantities

                                      = 1/2* (360-380)(245-330) = 850